 The synthesis of urea, [(NH2)2CO, is considered to be the first recognized example of preparing a biological compound from nonbiological reactants, challenging the notion that biological processes involved a “vital force” present only in living systems. Urea is also used in fertilizers of the plants. Today urea is produced industrially by reacting ammonia with carbon dioxide:  

2NH3(g) + CO2(g) → (NH2)2CO(aq) + H2O(l) 

 

In one process, 637.2 g of NH3 are treated with 1142 g of CO2.  

 

Which of the two reactants is the limiting reagent?  

 

 

Calculate the mass of (NH2)2CO formed.  

 

 

How much excess reagent (in grams) is left at the end of the reaction? 



1
Expert's answer
2021-02-09T03:47:20-0500
"n(NH3)=m\/M=637.2\/17=37.5 mol"

"n(CO2)=1142\/44=25.9 mol"

The limiting reagent is NH3. According to the reaction from 37.5 mol NH3 we can get 37.5/2=18.75 mol of (NH2)2CO.


"m((NH2)2CO)=n*M=18.75*60=1125 g"

25.9 - 18.75 = 7.15 mol CO2 is left at the end of the reaction.


"m(CO2)=7.15*44=314.6 g"